Brandy12:

The lengths of two sides of a triangle are shown. Side 1: 8x2 − 5x − 2 Side 2: 7x − x2 + 3 The perimeter of the triangle is 4x3 − 3x2 + 2x − 6. Part A: What is the total length of the two sides, 1 and 2, of the triangle? Show your work. (4 points) Part B: What is the length of the third side of the triangle? Show your work. (4 points) Part C: Do the answers for Part A and Part B show that the polynomials are closed under addition and subtraction? Justify your answer. (2 points)

yourdootcome:

part a = side 1 + side 2 = 7x^2 +2x +1 part b = perimeter - side a - side b = 4x^3 -10x^2 -7

jhonyy9:

@yourdootcome wrote:
part a = side 1 + side 2 = 7x^2 +2x +1 part b = perimeter - side a - side b = 4x^3 -10x^2 -7
please explain your calcules (work) !!!

jhonyy9:

@yourdootcome wrote:
part a = side 1 + side 2 = 7x^2 +2x +1 part b = perimeter - side a - side b = 4x^3 -10x^2 -7
...and what about part C. ?

jhonyy9:

@brandy12 do you understand these above wrote by @yourdootcome ?